OpenStudy (anonymous):

derivative of y=ln square root (1+sinx/1-sinx)

OpenStudy (anonymous):

ok let \[u = \frac{ 1+\sin(x) }{ 1-\sin(x) }\]

OpenStudy (anonymous):

use the Quotient rule\[\frac{ vdu-udv }{ v^2 }\]

OpenStudy (anonymous):

Better be specific though... \[\huge \frac{d}{dx}\frac{f(x)}{g(x)}=\frac{f'(x)g(x)-f(x)g'(x)}{[g(x)]^2}\]

OpenStudy (anonymous):

\[\frac{ (1-\sin(x))(\cos(x))-(1+\sin(x))(\cos(x)) }{ (1-\sin(x))^2 }\]

OpenStudy (anonymous):

simplify the above expression

OpenStudy (anonymous):

\[y =\frac{ 1 }{ 2 }\ln(u)\]

OpenStudy (anonymous):

\[\frac{ dy }{ du }=\frac{ 1 }{ 2 }\frac{ 1 }{ u }\]

OpenStudy (anonymous):

apply the chain rule

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ dy }{ du}\frac{ du }{ dx }\]

OpenStudy (anonymous):

@mathsmind Something's not quite right... I think it should be \[\large\frac{ (1-\sin(x))(\cos(x))-(1+\sin(x))(-\cos(x)) }{ (1-\sin(x))^2 }\]

OpenStudy (anonymous):

yes the minus sign i deleted by mistake i just noticed

OpenStudy (anonymous):

final answer is

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }\frac{ \ln(\frac{ \cos(x) }{ 1-\sin(x) }+\frac{ (1+\sin(x))\cos(x) }{ (1-\sin(x))^2 }) }{ \sqrt{\frac{ 1+\sin(x) }{ 1-\sin(x) }} }\]

OpenStudy (anonymous):

recall that \[\frac{ d }{ dx }\sin(x) = \cos(x)\]

OpenStudy (anonymous):

A bit hard to read, @mathsmind \[\huge \frac{ 1 }{ 2 }\frac{ \ln(\frac{ \cos(x) }{ 1-\sin(x) }+\frac{ (1+\sin(x))\cos(x) }{ (1-\sin(x))^2 }) }{ \sqrt{\frac{ 1+\sin(x) }{ 1-\sin(x) }} }\]

OpenStudy (anonymous):

\[\frac{ d }{ dx}(1+\sin(x))=\cos(x)\]

OpenStudy (anonymous):

change the resolution of the desktop or use opera browser

OpenStudy (anonymous):

or just place \large or \huge before the entire thingy... :)

OpenStudy (anonymous):

hehehe whatever

OpenStudy (anonymous):

You couldn't be bothered to type an extra five characters? -.-

OpenStudy (anonymous):

i am trying to get used to latex, so all my mind is focusing on coding

OpenStudy (anonymous):

i need to memorize the user manual better

OpenStudy (anonymous):

your resolution is too low that's why you are getting big fonts, at least set windows to small fonts

OpenStudy (anonymous):

from the control pannel

OpenStudy (anonymous):

Too much trouble :D

OpenStudy (anonymous):

hehehe ok

OpenStudy (anonymous):

i can see the full equation

OpenStudy (anonymous):

acids taste sour and bases taste bitter

OpenStudy (anonymous):

Where does this come from? 0.o

OpenStudy (anonymous):

how did this post come into mathematics? errrr